# arXiv:2603.00269v1[stat.ME] 27 Feb 2026

### v1

@@ Abstract
Linear regression estimators are known to be sensitive to outliers, and one alternative to obtain a robust and efficient estimator of the regression parameter is to model the error with Student’s $t$ distribution. In this article, we compare estimators of the degrees of freedom parameter in the $t$ distribution using frequentist and Bayesian methods, and then study properties of the corresponding estimated regression coefficient. We also include the comparison with some recommended approaches in the literature, including fixing the degrees of freedom and robust regression using the Huber loss. Our extensive simulations on both synthetic and real data demonstrate that estimating the degrees of freedom via the adjusted profile log-likelihood approach yields regression coefficient estimators with high accuracy, performing comparably to the maximum likelihood estimator where the degrees of freedom are fixed at their true values. These findings provide a detailed synthesis of $t$-based robust regression and underscore a key insight: the proper calibration of the degree of freedom is as crucial as the choice of the robust distribution itself for achieving optimal performance.
@@ 1 Introduction
In this paper we consider robust estimation in the linear regression model
@@
where $y=(y_{1},\dots,y_{n})$ is a vector of responses, $X$ is an $n\times p$ matrix of covariates with $i^{th}$ row $x_{i}^{T}$, and $\epsilon=(\epsilon_{1},\dots,\epsilon_{n})$ are unobserved random terms of mean zero and variance 1, independent of $y$ and $X$. Following several authors (Lange et al., 1989; Zellner, 1976; Singh, 1988; Fernández and Steel, 1999), we assume $\epsilon_{i},i=1,\dots,n$ are independent and identically distributed as Student’s $t$ with $\nu$ degrees of freedom.
Under this assumption, the likelihood function for (1) is
@@
Whether the covariates $X$ are fixed or random, inference is based on the conditional density of $y$ given $X$, as is usual in regression problems. The estimates derived from (2) may be expected to be more robust than least-squares estimates, as the $t$-distribution has heavier tails than the normal distribution, especially for smaller values of $\nu$; see Figure 1. As $\nu\rightarrow\infty$, the $t$-density function converges to the normal density function.
The ordinary least squares estimator of $\beta$ has maximum efficiency under the normal model, but is also sensitive to outliers. Models with smaller values of $\nu$ can accommodate heavier tails, but may be inefficient if the data suggests a larger value of $\nu$. Lange et al. (1989) suggested estimating $\nu$ when $n$ is large but choosing some fixed value if $n$ is small. Liu and Rubin (1995) described an extended expectation-maximization (EM) algorithm for computing maximum likelihood estimates of the $(\beta,\sigma,\nu)$. Singh (1988) proposed a moment-based estimator of $\nu$ for $\nu>4$. Several authors have considered Bayesian inference for $\nu$, using various choices of noninformative priors, such as Jeffreys’ prior (Fonseca et al., 2008; Fernández and Steel, 1999), a reference prior (He et al., 2021) and a uniform prior (Relles and Rogers, 1977).
Figure 1: Student’s ttdistribution has heavier tails than the normal distribution, with the difference decreasing as $\nu$ increases. Asymptotically, as $\nu\rightarrow\infty$ the $t$-distribution converges to the normal distribution.
In Section 2 we compare frequentist and Bayesian approaches to the estimation of $\nu$. In Section 3 we focus on inference for $\beta$, using a two-stage procedure in which we first estimate the degrees of freedom parameter $\nu$, and then fix $\nu$ at this estimated value in the likelihood function and maximize over $\beta$ and $\sigma$. With large $p$ we find that estimating $\nu$ by maximizing the adjusted profile likelihood function gives mean-squared error for $\beta$ comparable to that from minimizing Huber loss Huber (1973). In Section 4 we consider the robustness of our proposed methods in the presence of outliers.
@@ 2.1 Likelihood Approaches
$\ell_{\rm p}(\nu;y)=\ell(\hat{\beta}_{\nu},\hat{\sigma}_{\nu},\nu;y),$ (3)
where $\ell(\theta)=\log L(\theta;y)$, and $(\widehat{\beta}_{\nu},\widehat{\sigma}_{\nu})=\arg\max_{(\beta,\sigma)}\ell(\beta,\sigma,\nu;y)$ are the constrained maximum likelihood estimates of $(\beta,\sigma)$. Maronna (1976) and Zellner (1976) showed that $\widehat{\beta}_{\nu}$ and $\widehat{\sigma}_{\nu}$ exist. The asymptotic properties of $\hat{\nu}$ follow from classical likelihood theory:
$\sqrt{n}\,j_{\rm p}^{1/2}(\hat{\nu})(\hat{\nu}-\nu)\xrightarrow{d}N(0,1),\quad n\rightarrow\infty,$
@@
where $j_{\lambda\lambda}(\nu,\widehat{\lambda}_{\nu})$ is the sub-matrix of the observed Fisher information matrix that corresponds to the nuisance parameters. An expression for $j_{\lambda\lambda}(\nu,\lambda)$ is given in the supplementary material. Our second likelihood-based estimate is $\hat{\nu}_{adj}$, the value that maximizes the adjusted profile log-likelihood function $\ell_{\rm adj}(\nu;y)$.
@@ 2.2 Bayesian Approaches
where the prior for $(\sigma,\nu)$ is the determinant of the $2\times 2$ submatrix of (5). Using these priors we maximized the posterior distribution over $(\beta,\sigma,\nu)$ and extracted the Bayesian MAP estimate $\hat{\nu}_{Bayes}$.
An alternative partially Bayes approach is to treat the profile likelihood function for $\nu$ as a true likelihood function and use the prior $\pi(\nu)\propto I_{\nu\nu}^{1/2}(\nu)$ to form a pseudo-posterior distribution. The resulting posterior distribution is proper, as proved in the supplementary material. The MAP estimate of $\nu$ obtained from maximizing this pseudo-posterior distribution is denoted by $\hat{\nu}_{Pseudo}$.
Estimates of $\nu$ from these objective functions are invariant to location-scale transformations. Specifically, for the linear model $y_{i}=x_{i}\beta+\epsilon_{i}$, transforming the response as $y^{\prime}_{i}=ay_{i}+x_{i}\vec{b}$ with $a>0$ and a constant vector $\vec{b}$ shifts the log likelihoods and log pseudo-posterior by a constant, leaving the $\nu$ optimization results unchanged. See the supplementary material for details.
@@ 2.3 Simulations
In this section, we compare the performance of the four $\nu$-estimation approaches proposed above. We generated data from (1) with $t$ errors for $\epsilon_{i}$, with degrees of freedom $\nu=\{2,5,10\}$, $\sigma=1$, and $\beta\in\mathbb{R}^{p}$ with dimension $p\in\{1,2,5,10,20,40,60,80\}$. The covariates were generated from a multivariate standard normal distribution $\mathbf{N}_{p}(\boldsymbol{0},\mathbf{I}_{p})$. The choice of the true values of $\beta$ and $\sigma$ is arbitrary as the inference for $\nu$ is invariant to location-scale changes. The sample sizes were 300, 2500 and 4500 when $\nu=2,5,$ and $10$ (regardless of dimension $p$) respectively.
The BFGS algorithm provided in R (R Core Team, 2023) was used for all optimization steps in our simulations. This is a quasi-Newton method which iteratively approximates the Hessian of the objective function. The starting values of $(\beta,\sigma)$ for constrained maximum likelihood estimation were the ordinary least squares estimate of $\beta$ and usual residual mean square estimate of $\sigma$.
All computations were carried out in the $\omega$ parameterization, with initial guess set to the true value of $1/\nu$, but results are reported for estimation of $\nu$. The covariates were fixed in each set of simulation, to reflect modelling the conditional expected value of $y$ given $x$. We ran $500$ simulations for each combination of $\nu$ and $p$. The supplementary material provides more details about the choice of optimization details, including starting values, sample sizes and handling out-of-bound values.
@@
Table 1 presents the root mean-square-error (RMSE) averaged over the simulations, for the four estimators of $\nu$, when the true value of $\nu=2$ and the sample size $n=300$. For small values of $p$ the maximum profile likelihood estimator $\hat{\nu}$ has the smallest RMSE, but all four methods are comparable. For $p=60$ or $80$, the adjusted profile likelihood estimator $\hat{\nu}_{adj}$ has the smallest RMSE, and seems to adapt well to high-dimensional settings. The Bayesian estimators have similar RMSE when compared to each other and to the likelihood estimators. Other choices of $\nu$, $n$, and $p$, reported in the supplementary material, give similar conclusions. If $p$ is not large compared to $n$, we recommend the $\hat{\nu}$ and for larger $p$, $\hat{\nu}_{adj}$.
3 Estimating the regression
@@ 3.1 Simulation Settings
In linear regression, estimating the coefficient is the primary objective. To evaluate the effect of $\nu$-estimation in supporting coefficient estimation, we designed a simulation study using the stack loss data of Brownlee (1965). It has 21 responses ($Y$) in an industrial process, along with three explanatory variables: air flow ($X_{1}$), temperature ($X_{2}$), and acid concentration ($X_{3}$). We conducted simulations according to the following three data generation methods:
@@
    In the second method, we expanded the data to $p=\{40,80,120\}$ and $n=210$ to study the performance in a higher-dimensional regime with a larger sample size. The first four predictors (including the intercept) were based on the original stackloss predictors, in which we retained the original 21 observations, and generated an additional 189 observations with 3 predictors from a multivariate normal distribution with the same empirical mean vector and covariance matrix obtained from the original design matrix. All the remaining p4p-4 predictors were generated independently from a standard normal distribution. The corresponding true $\beta$ vector had the first four coefficients being the OLS estimates from the original data, and the remaining coefficients set as zero.
@@
We considered three approaches for the estimation of $\nu$: profile likelihood $\hat{\nu}_{p}$, adjusted profile likelihood $\hat{\nu}_{adj}$, and Jeffreys $\hat{\nu}_{Bayes}$. We also considered some fixed value other than the true $\nu=2$. For Bayesian methods, we only considered the Jeffreys approach because both Bayesian methods give very similar $\nu$-estimation in Section 2, and full independence Jeffreys is easier to implement, as it does not first construct the profile likelihood function. In each simulation, we estimated or fixed $\nu$ from the four approaches, then obtained $(\widehat{\beta}_{\nu},\widehat{\sigma}_{\nu})$ by optimizing the Student’s $t$ likelihood function. We also considered the ordinary least squares (OLS) and Huber regression estimators (Huber, 1973) of $\beta$ for comparison. For Huber’s method, the tuning parameter was automatically selected using the rlmDataDriven R package (Wang et al., 2019). This procedure was repeated over 500 simulations, again keeping the design matrix $X$ fixed while varying only the error term.
OLS imposes the largest penalty on the outliers compared to other methods. $t_{1}$ imposes lighter penalty than Huber regression with tuning parameter $c=1.5$). As $\nu$ increases, the loss function behaves more similarly as OLS loss function. See Figure 2.
@@
The simulation RMSE for $\hat{\beta}$ was computed as the average of the square root MSE over the scalar, ignoring the intercept
$RMSE(\hat{\beta})=\frac{1}{500}\sum_{k=1}^{500}\sqrt{\frac{1}{p-1}\sum_{j=2}^{p}(\beta_{jk}-\hat{\beta}_{jk})^{2}}.$ (8)
where $\hat{\beta}_{jk}$ is the $k$th simulation’s estimate of $j$th coefficient.
@@ 3.2 Results
The plots compare fixed $\nu$ approaches (in blue) for $\nu\in\{1,\dots,10\}\cup\{20,30\}$ and other traditional estimators (OLS and Huber) with the proposed $\nu$-estimation approaches (in red).
From Figure 3(a), we observe that the lowest overall RMSE is achieved when fixing $\nu=3$, with the RMSE at the true value $\nu=2$ is only slightly higher, indicating that the method is not highly sensitive to small misspecifications of $\nu$ around the true value. The slightly better performance at $\nu=3$ compared to the true $\nu=2$ is likely due to sampling variability in the simulation. As $n$ increases, the true $\nu$ would result in the lowest RMSE. For higher values of $\nu$ (e.g., $\nu=20$ or $\nu=30$), the RMSE for the estimate of $\beta$ increases notably. The three $\nu$-estimation approaches are comparable to the best fixed-$\nu$ results, and all outperform ordinary least squares. Estimation of $\beta$ using $\hat{\nu}_{adj}$ gives the smallest RMSE among $\nu$ estimated methods, which is comparable to RMSE of fixed $\nu=6$. Huber regression is slightly better than $\hat{\nu}_{adj}$ approach.
According to Figures 3(b) and 3(c), when $p=40,80$ ($p/n\approx 0.190,0.381$ respectively), the lowest overall RMSE is achieved by fixing $\nu$ at $3$ or $4$, and the RMSE increases for $\nu>4$. The $\hat{\nu}_{adj}$ approach outperforms Huber regression and gives the lowest RMSE among all $\nu$ estimated methods, which is comparable to fixing at the true $\nu$. In Figure 3(d), as we further increase $p$ to $120$ ($p/n\approx 0.571$), the lowest overall RMSE is achieved at larger fixed $\nu=7$. High dimensionality of $\beta$ gives the model enough flexibility to absorb extreme values into the fitted coefficients, shrinking the residuals relative to the true $t_{2}$ errors. With these outliers muted, the likelihood favours a larger degrees of freedom $\nu$, similar to selecting a less robust (steeper) loss function in Figure2, since the data no longer exhibit the heavy-tailed behaviour that would justify a small $\nu$. Among the $\nu$-estimation approaches, $\hat{\nu}_{adj}$ approach performs the best, its RMSE value is relatively large compared to the lowest RMSE that could be achieved by traditional methods. In particular, Huber regression performs better than the $\hat{\nu}_{adj}$ approach.
@@
The plots compare non $\nu$-estimation approaches (in blue) across various fixed values of $\nu$ and other traditional estimators (OLS and Huber) with the proposed $\nu$-estimation approaches (in red) in terms of overall RMSE.
According to Figures 4(a) and 4(b), the trend in overall RMSE using the simulated data with t error is similar to that of the stack loss data when $p/n<0.5$, where RMSE tends to increase as the fixed $\nu$ shifts away from the true value. The $\hat{\nu}_{adj}$ approach results in a lower RMSE than Huber regression. According to Figure 4(c), when $p/n=0.5$, we observe a similar RMSE trend as the stackloss $n=210,p=120$ simulations. The lowest RMSE is achieved at fixed $\nu=6$, which is not close to true $\nu=2$. Huber regression performs better than the $\hat{\nu}_{adj}$ approach, although both approaches give relatively large RMSEs compared to the smallest RMSE given by fixing $\nu=6$.
@@
The plot compares non $\nu$-estimation approaches (in blue) across various fixed values of $\nu$ and other traditional estimators (OLS and Huber) with the proposed $\nu$-estimation approaches (in red) in terms of overall RMSE.
With standard normal errors (Figures 5(a) and 5(b)), OLS performs the best in both $n=20$ and $n=500$ cases. All $\nu$-estimation approaches outperform Huber regression with $\hat{\nu}_{adj}$ approach coming closest to OLS among the $\nu$-estimation approaches. In the high-dimensional case $n=100;p=50$, Figure 5(c) shows that OLS again performs the best and only $\hat{\nu}_{adj}$ surpasses Huber regression among the $\nu$-estimation approaches. Overall, estimating $\nu$ is beneficial when when $p/n$ is not too large, as it would give promising results regardless the underlying error distribution. For errors with light-tail distribution, adjusted profile likelihood approach is comparable to the OLS, and for heavy tail distribution, it outperforms Huber when $p/n$ ratio is below 0.5. Although Lange et al. (1989) recommends fixing $\nu$ for small sample size $n$, the adjusted profile likelihood approach yields performance comparable to the best fixed-$\nu$ choices across error settings, supporting its use for estimating $\nu$.

### v3

# Robust Regression with Student’s $T$: The Role of Degrees of Freedom

@@ Abstract
Linear regression estimators are known to be sensitive to outliers, and one alternative to obtain a robust and efficient estimator of the regression parameter is to model the error with Student’s $t$ distribution. In this article, we compare estimators of the degrees of freedom parameter in the $t$ distribution using frequentist and Bayesian methods, and then study properties of the corresponding estimated regression coefficient. We also include the comparison with some recommended approaches in the literature, including fixing the degrees of freedom and robust regression using the Huber loss. Our extensive simulations on both synthetic and real data demonstrate that estimating the degrees of freedom via the adjusted profile log-likelihood approach yields regression coefficient estimators with high accuracy, performing comparably to the maximum likelihood estimators where the degrees of freedom are fixed at their true values. These findings provide a detailed synthesis of $t$-based robust regression and underscore a key insight: the proper calibration of the degrees of freedom is as crucial as the choice of the robust distribution itself for achieving optimal performance. The R package that implements our method is available at RobustTRegression.
@@ 1 Introduction
In this paper, we consider robust estimation in the linear regression model
@@
where $y=(y_{1},\dots,y_{n})$ is a vector of responses, $X$ is an $n\times p$ matrix of covariates with the $i$th row $x_{i}^{T}$, and $\epsilon=(\epsilon_{1},\dots,\epsilon_{n})$ are unobserved random terms of mean zero and variance 1, independent of $y$ and $X$. Following several studies (Lange et al., 1989; Zellner, 1976; Singh, 1988; Fernández and Steel, 1999), we assume $\epsilon_{i},i=1,\dots,n$ are independent and identically distributed as Student’s $t$ with $\nu$ degrees of freedom.
Under this assumption, the likelihood function under (1) is
@@
Regardless of whether the covariates $X$ are treated as fixed or random, inference is based on the conditional density of $y$ given $X$, as is usual in regression problems. The estimates derived from (2) may be expected to be more robust than least-squares estimates, as the $t$-distribution has heavier tails than the normal distribution, especially for smaller values of $\nu$; see Figure 1. As the degrees of freedom $\nu\rightarrow\infty$, the corresponding $t$-density function converges to the normal density function.
The ordinary least squares estimator of $\beta$ is most efficient under the normal model but is also sensitive to outliers. Student’s $t$ models with smaller values of $\nu$ can accommodate heavier tails, but may be inefficient if the data suggests a larger value of $\nu$. Lange et al. (1989) suggested estimating $\nu$ when $n$ is large but choosing some fixed value if $n$ is small. Liu and Rubin (1995) described an extended expectation-maximization (EM) algorithm for computing maximum likelihood estimates of the $(\beta,\sigma,\nu)$. Singh (1988) proposed a moment-based estimator of $\nu$ for $\nu>4$. Several studies have considered Bayesian inference for $\nu$, using various choices of noninformative priors, such as Jeffreys’ prior (Fonseca et al., 2008; Fernández and Steel, 1999), a reference prior (He et al., 2021), and a uniform prior (Relles and Rogers, 1977).
Figure 1: Student’s $t$ distribution has heavier tails than the normal distribution, with the difference decreasing as $\nu$ increases. Asymptotically, as $\nu\rightarrow\infty$ the $t$-distribution converges to the normal distribution.
In Section 2, we compare frequentist and Bayesian approaches to the estimation of $\nu$. In Section 3, we focus on inference for $\beta$ using a two-stage procedure, in which we first estimate the degrees of freedom parameter $\nu$, and then fix $\nu$ at this estimated value in the likelihood function and maximize over $\beta$ and $\sigma$. With large $p$, we find that estimating $\nu$ by maximizing the adjusted profile likelihood function gives mean squared error for $\beta$ comparable to that from minimizing the Huber loss (Huber, 1973). In Section 4, we consider the robustness of our proposed method in the presence of outliers. For reproducibility, the R (R Core Team, 2023) package for the implementations of our method and some competing methods considered in this paper are publicly available on GitHub at RobustTRegression.
@@ 2.1 Likelihood-Based Approaches
$\ell_{\rm p}(\nu;y)=\ell(\hat{\beta}_{\nu},\hat{\sigma}_{\nu},\nu;y):$ (3)
where $\ell(\theta)=\log L(\theta;y)$, and $(\widehat{\beta}_{\nu},\widehat{\sigma}_{\nu})=\arg\max_{(\beta,\sigma)}\ell(\beta,\sigma,\nu;y)$ are the constrained maximum likelihood estimators of $(\beta,\sigma)$. Maronna (1976) and Zellner (1976) showed that $\widehat{\beta}_{\nu}$ and $\widehat{\sigma}_{\nu}$ exist. The asymptotic properties of $\hat{\nu}$ follow from classical likelihood theory:
$j_{\rm p}^{1/2}(\hat{\nu})(\hat{\nu}-\nu)\xrightarrow{d}N(0,1),\quad n\rightarrow\infty,$
@@
where $j_{\lambda\lambda}(\nu,\widehat{\lambda}_{\nu})$ is the submatrix of the observed Fisher information matrix that corresponds to the nuisance parameters. An expression for $j_{\lambda\lambda}(\nu,\lambda)$ is given in the supplementary material. We propose a second likelihood-based estimator, $\hat{\nu}_{adj}$, defined as the value that maximizes the adjusted profile log-likelihood function $\ell_{\rm adj}(\nu;y)$.
@@ 2.2 Bayesian Approaches
where the prior for $(\sigma,\nu)$ is the determinant of the $2\times 2$ submatrix of (5). Using these priors, we maximized the posterior distribution over $(\beta,\sigma,\nu)$ and extracted the Bayesian maximum a posteriori (MAP) estimate $\hat{\nu}_{Bayes}$.
An alternative partially Bayesian approach is to treat the profile likelihood function for $\nu$ as a true likelihood function and use the prior $\pi(\nu)\propto I_{\nu\nu}^{1/2}(\nu)$ to form a pseudo-posterior distribution. The resulting posterior distribution is proper, as proved in the supplementary material. The MAP estimate of $\nu$ obtained from maximizing this pseudo-posterior distribution is denoted by $\hat{\nu}_{Pseudo}$.
Estimates of $\nu$ from these objective functions are invariant to location-scale transformations. Specifically, for the linear model $y_{i}=x_{i}\beta+\epsilon_{i}$, transforming the response as $y^{\prime}_{i}=ay_{i}+x_{i}b$ with $a>0$ and a constant vector $b$ shifts the log likelihoods and log pseudo-posterior by a constant, leaving the $\nu$ optimization results unchanged. See the supplementary material for details.
@@ 2.3 Simulations
In this section, we compare the performance of the four $\nu$-estimation approaches described above. We generated data from (1) with $t$ errors for $\epsilon_{i}$, with degrees of freedom $\nu\in\{2,5,10\}$, $\sigma=1$, and $\beta\in\mathbb{R}^{p}$ with dimension $p\in\{1,2,5,10,20,40,60,80\}$. The covariates were generated from a multivariate standard normal distribution $\mathbf{N}_{p}(\boldsymbol{0},\mathbf{I}_{p})$. The choice of the true values of $\beta$ and $\sigma$ is arbitrary, as the inference for $\nu$ is invariant to location-scale changes. The sample sizes were 300, 2500 and 4500 when $\nu=2,5,$ and $10$ (regardless of dimension $p$) respectively.
The BFGS algorithm provided in R (R Core Team, 2023) was used for all optimization steps in our simulations. This is a quasi-Newton method which iteratively approximates the Hessian of the objective function. The starting values of $(\beta,\sigma)$ for constrained maximum likelihood estimation were the ordinary least squares estimate of $\beta$ and the usual residual mean square estimate of $\sigma$.
All computations were carried out under the $\omega$ parameterization, with an initial guess set to the true value of $1/\nu$, but results are reported for estimation of $\nu$. The covariates were fixed in each set of simulations to reflect modelling the conditional expected value of $y$ given $x$. We ran $500$ simulations for each combination of $\nu$ and $p$. The supplementary material provides more details about the choice of optimization details, including starting values, sample sizes and handling out-of-bound values.
@@
Table 1 presents the root mean square error (RMSE) averaged over the simulations, for the four estimators of $\nu$, when the true value of $\nu=2$ and the sample size $n=300$. For small values of $p$, the maximum profile likelihood estimator $\hat{\nu}$ has the smallest RMSE, but all four methods are comparable. For $p=60$ or $80$, the adjusted profile likelihood estimator $\hat{\nu}_{adj}$ has the smallest RMSE, and seems to adapt well to high-dimensional settings. The Bayesian estimators have similar RMSE when compared to each other and to the likelihood estimators. Other choices of $\nu$, $n$, and $p$, reported in the supplementary material, give similar conclusions. If $p$ is not large compared to $n$, we recommend the $\hat{\nu}$ and for larger $p$, $\hat{\nu}_{adj}$.
3 Estimating the regression coefficient
In linear regression, estimating the coefficient is the primary objective. To evaluate the effect of $\nu$-estimation in supporting coefficient estimation, we designed a simulation study using the stack loss data of Brownlee (1965). It has 21 responses ($Y$) in an industrial process, along with three explanatory variables: air flow ($X_{1}$), temperature ($X_{2}$), and acid concentration ($X_{3}$).
@@ 3.1 Simulation Settings
We conducted simulations according to the following three data generation methods:
@@
    In the second method, we expanded the data to $p\in\{40,80,120\}$ and $n=210$ to study the performance in a higher-dimensional regime with a larger sample size. The first four predictors (including the intercept) were based on the original stackloss predictors, in which we retained the original 21 observations and generated an additional 189 observations with 3 predictors from a multivariate normal distribution with the same empirical mean vector and covariance matrix obtained from the original design matrix. All the remaining p4p-4 predictors were generated independently from a standard normal distribution. The corresponding true $\beta$ vector had the first four coefficients being the OLS estimates from the original data and the remaining coefficients set as zero.
@@
We considered three approaches for the estimation of $\nu$: profile likelihood $\hat{\nu}$, adjusted profile likelihood $\hat{\nu}_{adj}$, and Jeffreys $\hat{\nu}_{Bayes}$. We also considered some fixed value other than the true $\nu=2$. For Bayesian methods, we only considered the Jeffreys approach because both Bayesian methods give very similar $\nu$-estimation in Section 2, and full independence Jeffreys is easier to implement, as it does not first construct the profile likelihood function. In each simulation, we estimated or fixed $\nu$ from the four approaches, then obtained $(\widehat{\beta}_{\nu},\widehat{\sigma}_{\nu})$ by optimizing the Student’s $t$ likelihood function. We also considered the ordinary least squares (OLS) and Huber regression estimators (Huber, 1973) of $\beta$ for comparison. For Huber’s method, the tuning parameter was automatically selected using the rlmDataDriven R package (Wang et al., 2019). This procedure was repeated over 500 simulations, again keeping the design matrix $X$ fixed while varying only the error term.
OLS imposes the largest penalty on the outliers compared to other methods. $t_{1}$ imposes a lighter penalty than Huber regression with tuning parameter $c=1.5$). As $\nu$ increases, the loss function behaves more similarly to the OLS loss function. See Figure 2.
@@
The simulation RMSE for $\hat{\beta}$ was computed across the 500 simulation repetitions, excluding the intercept term in the regression coefficient:
$RMSE(\hat{\beta})=\frac{1}{500}\sum_{k=1}^{500}\sqrt{\frac{1}{p-1}\sum_{j=2}^{p}(\hat{\beta}_{jk}-\beta_{jk})^{2}},$ (8)
where $\hat{\beta}_{jk}$ is the $k$th simulation’s estimate of the $j$th component of the regression coefficient.
@@ 3.2 Results
The plots compare fixed $\nu$ approaches (in blue) for $\nu\in\{1,\dots,10\}\cup\{20,30\}$ and other traditional estimators (OLS and Huber) with the $\nu$-estimation approaches considered (in red).
From Figure 3(a), we observe that the lowest overall RMSE is achieved when fixing $\nu=3$, with the RMSE at the true value $\nu=2$ is only slightly higher, indicating that the method is not highly sensitive to small misspecifications of $\nu$ around the true value. The slightly better performance at $\nu=3$ compared to the true $\nu=2$ is likely due to sampling variability in the simulation. As $n$ increases, the true $\nu$ would result in the lowest RMSE. For higher values of $\nu$ (e.g., $\nu=20$ or $\nu=30$), the RMSE for the estimate of $\beta$ increases notably. The three $\nu$-estimation approaches are comparable to the best fixed-$\nu$ results, and all outperform ordinary least squares. Estimation of $\beta$ using $\hat{\nu}_{adj}$ gives the smallest RMSE among $\nu$ estimated methods, which is comparable to the RMSE of fixed $\nu=6$. Huber regression is slightly better than the $\hat{\nu}_{adj}$ approach.
According to Figures 3(b) and 3(c), when $p=40,80$ ($p/n=0.190,0.381$ respectively), the lowest overall RMSE is achieved by fixing $\nu$ at $3$ or $4$, and the RMSE increases for $\nu>4$. The $\hat{\nu}_{adj}$ approach outperforms Huber regression and gives the lowest RMSE among all $\nu$ estimated methods, which is comparable to fixing at the true $\nu$. In Figure 3(d), as we further increase $p$ to $120$ ($p/n=0.571$), the lowest overall RMSE is achieved at larger fixed $\nu=7$. High dimensionality of $\beta$ gives the model enough flexibility to absorb extreme values into the fitted coefficients, shrinking the residuals relative to the true $t_{2}$ errors. With these outliers muted, the likelihood favours a larger value of the degrees of freedom $\nu$, similar to selecting a less robust (steeper) loss function in Figure 2, since the data no longer exhibit the heavy-tailed behaviour that would justify a small $\nu$. Among the $\nu$-estimation approaches, the $\hat{\nu}_{adj}$ approach performs the best; its RMSE value is relatively large compared to the lowest RMSE that could be achieved by traditional methods. In particular, Huber regression performs better than the $\hat{\nu}_{adj}$ approach.
@@
The plots compare non $\nu$-estimation approaches (in blue) across various fixed values of $\nu$ and other traditional estimators (OLS and Huber) with the $\nu$-estimation approaches considered (in red) in terms of overall RMSE.
According to Figures 4(a) and 4(b), the trend in the overall RMSE using the simulated data with $t$ error is similar to that of the stack loss data when $p/n<0.5$, where the RMSE tends to increase as the fixed $\nu$ shifts away from the true value. The $\hat{\nu}_{adj}$ approach results in a lower RMSE than Huber regression. According to Figure 4(c), when $p/n=0.5$, we observe a similar RMSE trend as the stackloss $n=210,p=120$ simulations. The lowest RMSE is achieved at fixed $\nu=6$, which is not close to the true $\nu=2$. Huber regression performs better than the $\hat{\nu}_{adj}$ approach, although both approaches give relatively large RMSEs compared to the smallest RMSE given by fixing $\nu=6$.
@@
The plot compares non $\nu$-estimation approaches (in blue) across various fixed values of $\nu$ and other traditional estimators (OLS and Huber) with the $\nu$-estimation approaches considered (in red) in terms of overall RMSE.
With standard normal errors (Figures 5(a) and 5(b)), OLS performs the best in both $n=20$ and $n=500$ cases. All $\nu$-estimation approaches outperform Huber regression, with the $\hat{\nu}_{adj}$ approach coming closest to OLS among the $\nu$-estimation approaches. In the high-dimensional case $n=100;p=50$, Figure 5(c) shows that OLS again performs the best, and only $\hat{\nu}_{adj}$ surpasses Huber regression among the $\nu$-estimation approaches. Overall, estimating $\nu$ is beneficial when $p/n$ is not too large, as it would give promising results regardless of the underlying error distribution. For errors with light-tailed distributions, the adjusted profile likelihood approach is comparable to the OLS, and for heavy-tailed distributions, it outperforms Huber when the $p/n$ ratio is below 0.5. Although Lange et al. (1989) recommends fixing $\nu$ for small sample size $n$, the adjusted profile likelihood approach yields performance comparable to the best fixed-$\nu$ choices across error settings, supporting its use for estimating $\nu$.